### v1

# A relative version of Bass’ theorem about finite-dimensional algebras

Abstract.
As a special case of Bass’ theory of perfect rings, one obtains the assertion that, over a finite-dimensional associative algebra over a field, all flat modules are projective. In this paper we prove the following relative version of this result. Let $R\longrightarrow A$ be a homomorphism of associative rings such that $A$ is a finitely generated projective right $R$-module. Then every flat left $A$-module is a direct summand of an $A$-module filtered by $A$-modules $A\otimes_{R}F$ induced from flat left $R$-modules $F$. In other words, a left $A$-module is cotorsion if and only if its underlying left $R$-module is cotorsion. The proof is based on the cotorsion periodicity theorem.
Introduction
All finite-dimensional associative algebras $A$ over a field $k$ are left and right Artinian rings. Any associative ring that is either left or right Artinian is left perfect. According to [1, Theorem P], it follows that all flat $A$-modules are projective.
This proof of projectivity of flat modules is based on the structure theory of finite-dimensional algebras over a field. Any such algebra $A$ has a radical $J$, which is a nilpotent ideal in $A$, and the quotient algebra $A/J$ is classically semisimple.
According to the orginal Bass’ definition in [1], an associative ring $A$ is said to be left perfect if all left $A$-modules have projective covers. By [1, Theorem P(3)], the ring $A$ is left perfect if and only if all flat left $A$-modules are projective. The condition of [1, Theorem P(1)] provides a characterization of left perfect rings in terms of their structural properties that is used in the argument above.
Now let $R$ be a commutative ring and $A$ be an associative $R$-algebra that is finitely generated and projective as an $R$-module. What can one say about flat $A$-modules in this relative context? Notice that the structure theory of finite-dimensional algebras over a field is not available for finite algebras over commutative rings.
Our approach to the question in the previous paragraph is based on the notion of a cotorsion pair, which was introduced by Salce [15] in the context of the theory of infinitely generated abelian groups. An introductory discussion of cotorsion pairs in abelian categories and module categories can be found in the introduction to the paper [12]. The book [9] contains a wealth of results and examples.
Let $A$ be an associative ring. A cotorsion pair $(\mathsf{F},\mathsf{C})$ in the category of left $A$-modules $A{\operatorname{\mathsf{--Mod}}}$ is a pair of classes of modules $\mathsf{F}$, $\mathsf{C}\subset A{\operatorname{\mathsf{--Mod}}}$ such that $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all $F\in\mathsf{F}$,  $C\in\mathsf{C}$, and both the classes $\mathsf{F}$ and $\mathsf{C}$ are maximal for this property with respect to each other.
The thematic example of a cotorsion pair if the flat cotorsion pair (also known as the Enochs cotorsion pair [8], [9, Definition 5.18(i)]). In this example, the class $\mathsf{F}=A{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ is the class of all flat left $A$-modules. The class $\mathsf{C}=A{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}$ is the class of all cotorsion left $A$-modules. Here a left $A$-module $C$ is said to be cotorsion if $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all flat left $A$-modules $F$, or equivalently, $\operatorname{Ext}^{n}_{A}(F,C)=0$ for all flat left $A$-modules $F$ and all $n\geq 1$.
The main result about cotorsion pairs is the Eklof–Trlifaj theorem [6, Theorems 2 and 10], [9, Theorem 6.11 and Corollary 6.14]. The Eklof–Trlifaj theorem concerns the notion of a complete cotorsion pair, and its simplest formulation claims that every cotorsion pair generated by a set of modules is complete. Proving completeness of the flat cotorsion pair in the category of modules over an arbitrary ring was a major open problem, which became known under the name of the Flat Cover Conjecture. Posed by Enochs in [7] in 1981 (see also [16, Section 1.3]), it was proved by Bican–El Bashir–Enochs in [3] in 2001. One of the two proofs of the Flat Cover Conjecture presented in [3] was based on the Eklof–Trlifaj theorem.
A trivial example of a cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$ is the projective cotorsion pair. It is formed by the two classes of projective $A$-modules $\mathsf{F}=A{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}}$ and arbitrary $A$-modules $\mathsf{C}=R{\operatorname{\mathsf{--Mod}}}$. The assertion that the projective cotorsion pair is complete means precisely that there are enough projective objects in $A{\operatorname{\mathsf{--Mod}}}$.
Another cotorsion pair relevant for the purposes of the present paper is the flaprojective cotorsion pair. It depends on a homomorphism of associative rings $R\longrightarrow A$. The flaprojective cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$ is a relative mixture of the flat cotorsion pair in $R{\operatorname{\mathsf{--Mod}}}$ and the projective cotorsion pair in the direction of $A$ relative to $R$.
Specifically, the flaprojective cotorsion pair $(\mathsf{F},\mathsf{C})$ is formed by the following two classes of left $A$-modules. The right-hand class $\mathsf{C}$ consists of all left $A$-modules that are cotorsion as $R$-modules, i. e., the $A$-modules whose underlying $R$-module is cotorsion. The left-hand class $\mathsf{F}$ consists of all the left $A$-modules $F$ such that $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion over $R$. The $A$-modules $F$ satisfying the latter condition are called flaprojective relative to $R$ (or $A/R$-flaprojective) [13, Section 1.6].
The definition of a flaprojective module in the previous paragraph is very indirect and roundabout. First one defines cotorsion $R$-modules as the $R$-modules right $\operatorname{Ext}^{1}_{R}$-orthogonal of flat $R$-modules, and then one defined flaprojective $A$-modules as the $A$-modules left $\operatorname{Ext}^{1}_{A}$-orthogonal to $R$-cotorsion $A$-modules. A much more direct and explicit description is provided by the Eklof–Trlifaj theorem, but it uses the concept of an ordinal-indexed filtration on modules. A left $A$-module $F$ is $A/R$-flaprojective if and only if $F$ is a direct summand of an $A$-module admitting an ordinal-indexed filtration whose successive quotients are the $A$-modules $A\otimes_{R}G$ induced from flat left $R$-modules $G$.
Now that we have explained the meaning of the concepts involved, we can formulate the main result of this paper.
Theorem 0.1.
Let $R\longrightarrow A$ be a homomorphism of associative rings making $A$ a finitely generated projective right $R$-module. Then the flat cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$ consides with the $A/R$-flaprojective cotorsion pair. More specifically:
(a) A left $A$-module $C$ is cotorsion if and only if $C$ is cotorsion as a left $R$-module.
(b) A left $A$-module is flat if and only if it is $A/R$-flaprojective. In other words, a left $A$-module $F$ is flat if and only if $F$ is a direct summand of an $A$-module filtered by $A$-modules $A\otimes_{R}G$ induced from flat left $R$-modules $G$.
It would be interesting to know how far one can generalize Theorem 0.1. Quite generally, for any homomorphism of associative rings $R\longrightarrow A$, property (a) in the theorem holds if and only if property (b) holds. If this is the case, one can say that the ring homomorphism $R\longrightarrow A$ is relatively left perfect.
In the absolute setting, a ring $A$ is called left perfect if all flat left $A$-modules are projective [1, Theorem P]. For example, if $A$ is an associative algebra over a field $k$, then the ring $A$ is left perfect in the sense of [1] if and only if the ring homomorphism $k\longrightarrow A$ is relatively left perfect in the sense of the previous paragraph.
There are many examples of left perfect rings and left perfect $k$-algebras beyond the finite-dimensional $k$-algebras. Are there any nontrivial classes of examples of relatively left perfect ring homomorphisms beyond the ones described by Theorem 0.1 ? Can one characterize such ring homomorphisms by any relative analogues of the properties listed in [1, Theorem P] ?
Acknowledgement
I am grateful to Michal Hrbek and Sergio Pavon for helpful conversations. The author is supported by the GAČR project 23-05148S and the Institute of Mathematics, Czech Academy of Sciences (research plan RVO: 67985840).
1. Preliminaries on Cotorsion Pairs
Let $A$ be an associative ring. We denote by $A{\operatorname{\mathsf{--Mod}}}$ the abelian category of left $A$-modules.
Let $\mathsf{F}$, $\mathsf{C}\subset A{\operatorname{\mathsf{--Mod}}}$ be two classes of $A$-modules. Then one denotes by $\mathsf{F}^{\perp_{1}}\subset A{\operatorname{\mathsf{--Mod}}}$ the class of all left $A$-modules $X$ such that $\operatorname{Ext}^{1}_{A}(F,X)=0$ for all $F\in\mathsf{F}$. Dually, the notation ${}^{\perp_{1}}\mathsf{C}\subset A{\operatorname{\mathsf{--Mod}}}$ stands for the class of all left $A$-modules $Y$ such that $\operatorname{Ext}^{1}_{A}(Y,C)=0$ for all $C\in\mathsf{C}$.
A pair of classes $(\mathsf{F},\mathsf{C})$ in $A{\operatorname{\mathsf{--Mod}}}$ is said to be a cotorsion pair if $\mathsf{C}=\mathsf{F}^{\perp_{1}}$ and $\mathsf{F}={}^{\perp_{1}}\mathsf{C}$. For any class of modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$, the pair of classes $\mathsf{F}=\mathsf{S}^{\perp_{1}}$ and $\mathsf{C}={}^{\perp_{1}}\mathsf{C}$ is a cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$. The latter cotorsion pair $(\mathsf{F},\mathsf{C})$ is said to be generated by the class of modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$.
A cotorsion pair $(\mathsf{F},\mathsf{C})$ in $A{\operatorname{\mathsf{--Mod}}}$ is said to be complete if for every left $A$-module $M$ there exist short exact sequences of $A$-modules
[…]
with $A$-modules $F$, $F^{\prime}\in\mathsf{F}$ and $C$, $C^{\prime}\in\mathsf{C}$. The short exact sequence (1) is called a special precover sequence. The short exact seqeuence (2) is called a special preenvelope sequence. Taken together, the short exact sequences (12) are called the approximation sequences.
An introductory discussion of cotorsion pairs can be found in the introduction to [12]. Many examples of approximation sequences (for abelian groups, i. e., modules over the ring of integers $A=\mathbb{Z}$) are discussed in [11, Section 12]. The main reference source on cotorsion pairs in module categories is the book [9].
Let $F$ be an $A$-module and $\alpha$ be an ordinal. An $\alpha$-indexed filtration on $F$ is a family of submodules $F_{i}\subset F$, indexed by the ordinals $0\leq i\leq\alpha$ and satisfying the following conditions:
  • $F_{0}=0$ and $F_{\alpha}=F$;
  • $F_{i}\subset F_{j}$ for all ordinals $0\leq i\leq j\leq\alpha$;
  • $F_{j}=\bigcup_{i<j}F_{i}$ for all limit ordinals $0<j\leq\alpha$.
Given an $A$-module $F$ with an ordinal-indexed filtration $(F_{i}\subset F)_{0\leq i\leq\alpha}$, one considers the successive quotient $A$-modules $S_{i}=F_{i+1}/F_{i}$, where $0\leq i<\alpha$. The $A$-module $F$ is said to be filtered by the $A$-modules $S_{i}$. In an alternative terminology, the $A$-module $F$ is said to be a transfinitely iterated extension (in the sense of the direct limit) of the $A$-modules $S_{i}=F_{i+1}/F_{i}$.
For any class of $A$-modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$, one denotes by $\mathsf{Fil}(\mathsf{S})$ the class of all $A$-modules filtered by (modules isomorphic to) modules from $\mathsf{S}$. A class of $A$-modules $\mathsf{F}\subset A{\operatorname{\mathsf{--Mod}}}$ is said to be deconstructible if there exists a set (rather than a proper class) of modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$ such that $\mathsf{F}=\mathsf{Fil}(\mathsf{S})$.
The following result is known as the Eklof lemma.
Lemma 1.1.
For any class of $A$-modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$, one has $\mathsf{S}^{\perp_{1}}=\mathsf{Fil}(\mathsf{S})^{\perp_{1}}$. In other words, for any class of $A$-modules $\mathsf{C}\subset A{\operatorname{\mathsf{--Mod}}}$, the class of $A$-modules ${}^{\perp_{1}}\mathsf{C}\subset A{\operatorname{\mathsf{--Mod}}}$ is closed under transfinitely iterated extensions, that is, ${}^{\perp_{1}}\mathsf{C}=\mathsf{Fil}({}^{\perp_{1}}\mathsf{C})$.
Proof.
This is [6, Lemma 1] or [9, Lemma 6.2]. ∎
Given a class of $A$-modules $\mathsf{F}\subset A{\operatorname{\mathsf{--Mod}}}$, we denote by $\mathsf{F}^{\oplus}\subset A{\operatorname{\mathsf{--Mod}}}$ the class of all direct summands of modules from $\mathsf{F}$.
Theorem 1.2 (Eklof–Trlifaj [6]).
Let $A$ be an associative ring and $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$ be a set of left $A$-modules. In this context:
(a) The cotorsion pair $(\mathsf{F},\mathsf{C})$ generated by $\mathsf{S}$ in $A{\operatorname{\mathsf{--Mod}}}$ is complete.
(b) Assume that the free left $A$-module $A$ with one generator belongs to $\mathsf{S}$, that is ${}_{A}A\in\mathsf{S}$. Then one has $\mathsf{F}=\mathsf{Fil}(\mathsf{S})^{\oplus}$.
Proof.
This is [6, Theorems 2 and 10] or [9, Theorem 6.11 and Corollary 6.14]. ∎
Corollary 1.3.
Let $A$ be an associative ring and $\mathsf{F}\subset A{\operatorname{\mathsf{--Mod}}}$ be a class of left $A$-modules. In this context:
(a) If the class $\mathsf{F}$ is deconstructible, then the cotorsion pair generated by $\mathsf{F}$ in $A{\operatorname{\mathsf{--Mod}}}$ is complete.
(b) If the class $\mathsf{F}$ is deconstructible, closed under direct summands in $A{\operatorname{\mathsf{--Mod}}}$, and contains the free $A$-module ${}_{A}A$, then $\mathsf{F}$ is the left-hand part of a complete cotorsion pair $(\mathsf{F},\mathsf{C})$ in $A{\operatorname{\mathsf{--Mod}}}$. In other words, the pair of classes $(\mathsf{F},\mathsf{F}^{\perp_{1}})$ is a complete cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$.
Proof.
Follows from Lemma 1.1 and Theorem 1.2. ∎
2. Cotorsion Modules
Let $R$ be an associative ring. A left $R$-module $C$ is said to be cotorsion (or Enochs cotorsion [8]) if $\operatorname{Ext}^{1}_{R}(F,C)=0$ for all flat left $R$-modules $F$.
Using the facts that all projective left $R$-modules are flat and the class of flat modules is closed under kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}$, one can easily show that $\operatorname{Ext}_{R}^{n}(F,C)=0$ for all flat left $R$-modules $F$, all cotorsion left $R$-modules $C$, and all integers $n\geq 1$ (cf. [9, Lemma 5.24] and [14, Lemma 7.1]). It follows that the class of cotorsion $R$-modules is closed under cokernels of monomorphisms in $R{\operatorname{\mathsf{--Mod}}}$; it is also obviously closed under extensions and infinite products. We denote the class of flat left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}\subset R{\operatorname{\mathsf{--Mod}}}$ and the class of cotorsion left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\subset R{\operatorname{\mathsf{--Mod}}}$.
The most nontrivial known property of cotorsion modules is the following cotorsion periodicity theorem of Bazzoni, Cortés-Izurdiaga, and Estrada [2].
Theorem 2.1.
Let $A$ be an associative ring and $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be an (unbounded) acyclic complex of $A$-modules such that all the $A$-modules $C^{n}$,  $n\in\mathbb{Z}$ are cotorsion. Then the $A$-modules of cocycles of the acyclic complex $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are cotorsion $A$-modules, too.
Proof.
This is [2, Theorem 5.1(2)]. See also [2, Theorem 1.2(2) or Proposition 4.8(2)] for essentially equivalent formulations of this theorem. A further discussion in category-theoretic context can be found in [14, Corollary 9.2 and Proposition 9.4]. ∎
Corollary 2.2.
Let $A$ be an associative ring and $M$ be an $A$-module. Assume that $M$ admist a (left) resolution by cotorsion $A$-modules, i. e., there exists an exact sequence of $A$-modules
[…]
with cotorsion $A$-modules $C_{n}$,  $n\geq 0$. Then $M$ is a cotorsion $A$-module.
Proof.
This is [10, Corollary 1.6.4]. The point is that any $A$-module can be embedded into a cotorsion $A$-module (in fact, all injective $A$-modules are cotorsion by definition). Therefore, any $A$-module $M$ has a coresolution (right resolution) by cotorsion $A$-modules, i. e., there is an exact complex of $A$-modules
[…]
with cotorsion $A$-modules $C^{n}$,  $n\geq 0$. Now the combined complex (the resulting two-sided resolution of $M$)
[…]
is an acyclic complex of cotorsion $A$-modules, and $M$ is one of its $A$-modules of cocycles. So Theorem 2.1 is applicable. ∎
The following three lemmas are more elementary. We use the notation $\operatorname{Hom}_{R}(L,M)$ for the abelian group of morphisms $L\longrightarrow M$ in the category of left $R$-modules $R{\operatorname{\mathsf{--Mod}}}$. The group of homomorphisms of right $R$-modules $N\longrightarrow Q$ will be denoted by $\operatorname{Hom}_{R^{\mathrm{op}}}(N,Q)$.
Lemma 2.3.
Let $R\longrightarrow A$ be a homomorphism of associative rings and $D$ be a cotorsion $A$-module. Then the underlying $R$-module of $D$ is a cotorsion $R$-module.
Proof.
This is [10, Lemma 1.3.4(a)] or [13, Lemma 1.6(a)]. One can use the $\operatorname{Ext}^{1}$-adjunction of change of rings as in [5, Proposition VI.4.1.3] (see Lemma 3.4 below). Alternatively, the following elementary argument works. A left $R$-module $C$ is cotorsion if and only if the functor $\operatorname{Hom}_{R}({-},C)$ takes every short exact sequence of flat left $R$-modules to a short exact sequence of abelian groups. Given a short exact sequence of flat left $R$-modules $0\longrightarrow H\longrightarrow G\longrightarrow F\longrightarrow 0$, we have a short exact sequence of flat left $A$-modules $0\longrightarrow A\otimes_{R}H\longrightarrow A\otimes_{R}G\longrightarrow A\otimes_{R}F\longrightarrow 0$. By assumption, the $A$-module $D$ is cotorsion, so the functor $\operatorname{Hom}_{A}({-},D)$ takes the latter short exact sequence of $A$-modules to a short exact sequence of abelian groups. It remain to recall the adjunction isomorphism of abelian groups $\operatorname{Hom}_{A}(A\otimes_{R}M,\>D)\simeq\operatorname{Hom}_{R}(M,D)$, which holds for any left $R$-module $M$. The assertion of the lemma is also a special case of the next Lemma 2.4. ∎
Lemma 2.4.
Let $R$ and $S$ be two associative rings, $K$ be an $R$-$S$-bimodule that is flat as a left $R$-module, and $C$ be a cotorsion left $R$-module. Then the left $S$-module $\operatorname{Hom}_{R}(K,C)$ is cotorsion.
Proof.
This is [10, Lemma 1.3.3(a)] or [13, Lemma 1.5(a)]. The point is that the tensor product $K\otimes_{S}G$ is a flat left $R$-module for any flat left $S$-module $G$. Based on this observation, one can use the $\operatorname{Ext}^{1}$-adjunction of the tensor product and Hom functors as in [5, formula (4) in Section XVI.4], [10, formula (1.1) in Section 1.2], or [13, Lemma 1.1]. Alternatively, an elementary argument can be found in [13, proof of Lemma 1.5]. ∎
Lemma 2.5.
Let $R\longrightarrow A$ be a homomorphism of associative rings making $A$ a finitely generated projective right $R$-module. Consider the $R$-$A$-bimodule
$K=\operatorname{Hom}_{R^{\mathrm{op}}}(A,R),$
with the left $R$-module structure on $K$ coming from the left $R$-module structure on $R$ and the right $A$-module structure on $K$ induced by the left $A$-module structure on $A$. Then, for every left $R$-module $M$, there is a natural isomorphism of left $A$-modules
$A\otimes_{R}M\simeq\operatorname{Hom}_{R}(K,M),$
where the left $A$-module structure on $A\otimes_{R}M$ comes from the left $A$-module structure on $A$ and the left $A$-module structure on $\operatorname{Hom}_{R}(K,M)$ comes from the right $A$-module structure on $K$.
Proof.
More generally, for any two associative rings $R$ and $S$, let $B$ be an $S$-$R$-bimodule that is finitely generated and projective as a right $R$-module, and let $M$ be a left $R$-module. Then the $R$-$S$-bimodule $L=\operatorname{Hom}_{R^{\mathrm{op}}}(B,R)$ is finitely generated and projective as a left $R$-module, there is a natural isomorphism of $S$-$R$-bimodules $B\simeq\operatorname{Hom}_{R}(L,R)$, and there is a natural isomorphism of left $S$-modules $B\otimes_{R}M\simeq\operatorname{Hom}_{R}(L,M)$. Now one can put $B=S=A$ and $L=K$. ∎
Proposition 2.6.
Let $R\longrightarrow A$ be a homomorphism of associative rings making $A$ a finitely generated projective right $R$-module. Let $C$ be a cotorsion left $R$-module. Then the left $A$-module $A\otimes_{R}C$ is cotorsion.
Proof.
By Lemma 2.5, we have a left $A$-module isomorphism $A\otimes_{R}C\simeq\operatorname{Hom}_{R}(K,C)$, where $K=\operatorname{Hom}_{R^{\mathrm{op}}}(A,R)$. The $R$-$A$-bimodule $K$ is finitely generated and projective, hence flat, as a left $R$-module. So Lemma 2.4 is applicable for the ring $S=A$. ∎
Lemma 2.7.
Let $R\longrightarrow A$ be a homomorphism of associative rings. Then every left $A$-module $M$ admits a natural relative bar resolution of the form
(3) $\dotsb\longrightarrow A\otimes_{R}A\otimes_{R}A\otimes_{R}M\longrightarrow A\otimes_{R}A\otimes_{R}M\longrightarrow A\otimes_{R}M\longrightarrow M\longrightarrow 0.$
Proof.
The differential in the relative bar resolution is given by the standard formula
$\partial(a_{0}\otimes a_{1}\otimes\dotsb\otimes a_{n}\otimes m)$ $=a_{0}a_{1}\otimes a_{2}\otimes a_{3}\otimes\dotsb\otimes a_{n}\otimes m$
$-a_{0}\otimes a_{1}a_{2}\otimes a_{3}\otimes a_{4}\otimes\dotsb\otimes a_{n}\otimes m+\dotsb$
$+(-1)^{n-1}a_{0}\otimes a_{1}\otimes\dotsb\otimes a_{n-2}\otimes a_{n-1}a_{n}\otimes m$
$+(-1)^{n}a_{0}\otimes a_{1}\otimes\dotsb\otimes a_{n-1}\otimes a_{n}m,$
for all $a_{0}$, …, $a_{n}\in A$,  $n\geq 0$, and $m\in M$. It is clear from this formula that $\partial$ is a left $A$-module morphism (where the left $A$-module structure on the tensor products comes from the left action of $A$ on the leftmost tensor factor $A$). One can easily compute that (3) is a complex, i. e., $\partial^{2}=0$.
The lemma claims that the complex (3) is acyclic. In fact, the complex (3) is contractible as a complex of left $R$-modules (but not as a complex of left $A$-modules). The natural $R$-linear contracting homotopy $h$ is given by the formula
$h(a_{1}\otimes a_{2}\otimes\dotsb\otimes a_{n}\otimes m)=1\otimes a_{1}\otimes a_{2}\otimes\dotsb\otimes a_{n}\otimes m,$
where $1\in A$ is the unit element. ∎
Now we can prove the main result of this section, which is also the first main result of this paper.
Theorem 2.8.
Let $R\longrightarrow A$ be a homomorphism of associative rings such that $A$ is a finitely generated projective right $R$-module. Then a left $A$-module $D$ is cotorsion as an $A$-module if and only if it is cotorsion as an $R$-module.
Proof.
The easy “only if” assertion holds by Lemma 2.3. To prove the “if”, consider the left $A$-module $A\otimes_{R}D$ (with the $A$-module structure coming from the left $A$-module structure on $A$). Since $D$ is a cotorsion left $R$-module by assumption, Proposition 2.6 tells us that $A\otimes_{R}D$ is a cotorsion left $A$-module.
By Lemma 2.3, it follows that left $R$-module $A\otimes_{R}D$ (with the $R$-module structure coming from the left $R$-module structure on $A$) is cotorsion. Therefore, Proposition 2.6 tells us that the left $A$-module $A\otimes_{R}A\otimes_{R}D$ (with the $A$-module structure coming from the left $A$-module structure on the leftmost tensor factor $A$) is cotorsion.
Proceeding in this way, we prove by induction that the left $A$-module $A^{\otimes_{R}\,n}\otimes_{R}D$ (with the $A$-module structure coming from the left $A$-module structure on the leftmost tensor factor $A$) is cotorsion for all $n\geq 1$. Now Lemma 2.7 provides a left resolution of $D$ by cotorsion $A$-modules. By Corollary 2.2, we can conclude that $D$ is a cotorsion $A$-module, as desired. ∎
3. Flaprojective Modules
Let $R$ be an associative ring. It is easy to prove that the pair of classes of flat left $R$-modules and cotorsion left $R$-modules $(\mathsf{F},\mathsf{C})$ $=$ $(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})$ is a cotorsion pair in $R{\operatorname{\mathsf{--Mod}}}$[16, Lemma 3.4.1], [13, Lemma 1.3]. Proving that this pair of classes, called the flat cotorsion pair or the Enochs cotorsion pair [8], [9, Definition 5.18(i)]), is a complete cotorsion pair, is harder. There are two proofs of this result in the paper [3], one of them based on the following observation of Enochs.
Proposition 3.1.
Let $R$ be an associative ring. Then the class of all flat left $R$-modules $\mathsf{F}=R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ is deconstructible in $R{\operatorname{\mathsf{--Mod}}}$. More precisely, let $\lambda$ be an infinite cardinal not smaller than the cardinality $|R|$ of the ring $R$, and let $\mathsf{S}$ be the set of (representatives of isomorphism classes of) flat left $R$-modules $S$ of the cardinality $|S|$ not exceeding $\lambda$. Then one has $\mathsf{F}=\mathsf{Fil}(\mathsf{S})$.
Proof.
This is [3, Lemma 1 and Proposition 2]. See [9, Lemma 6.17] for a generalization. ∎
Corollary 3.2.
For any associative ring $R$, the flat cotorsion pair $(\mathsf{F},\mathsf{C})$ $=$ $(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})$ in $R{\operatorname{\mathsf{--Mod}}}$ is complete.
Proof.
This is [3, Section 2] and a particular case of [9, Theorem 6.19]. The assertion follows from Proposition 3.1 by virtue of Corollary 1.3(b). ∎
Now let $R\longrightarrow A$ be a homomorphism of associative rings. A left $A$-module $F$ is said to be flaprojective relative to $R$ (or $A/R$-flaprojective) [13, Section 1.6] if $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion as left $R$-modules (i. e., the underlying left $R$-module of $C$ is cotorsion).
Since the class of cotorsion $R$-modules is closed under cokernels of monomorphisms (see the beginning of Section 2), so is the class of all $A$-modules that are cotorsion over $R$. All cotorsion $A$-modules are cotorsion over $R$ by Lemma 2.3; in particular, all injective $A$-modules are cotorsion over $R$.
Using these observations, one can show that $\operatorname{Ext}^{n}_{A}(F,C)=0$ for all $A/R$-flaprojective left $R$-modules $F$, all left $A$-modules $C$ that are cotorsion over $R$, and all integers $n\geq 1$ (cf. [9, Lemma 5.24] and [14, Lemma 7.1]). It follows that the class of $A/R$-flaprojective $A$-modules is closed under kernels of epimorphisms in $A{\operatorname{\mathsf{--Mod}}}$; it is also obviously closed under extensions and infinite direct sums. By Lemma 1.1, the class of $A/R$-flaprojective left $A$-modules is closed under transfinitely iterated extensions.
Remarks 3.3.
(1) Since all cotorsion $A$-modules are cotorsion over $R$, it follows that all $A/R$-flaprojective $A$-modules are flat [13, Remark 1.28]. Obviously, all projective $A$-modules are $A/R$-flaprojective; so the class of $A/R$-flaprojective $A$-modules is, generally speaking, intermediate between the classes of projective and flat $A$-modules.
(2) If $A=R$ and $R\longrightarrow A$ is the identity ring homomoprhism, then the class of $A/R$-flaprojective $A$-modules coincides with the class of flat $R$-modules [13, Example 1.27(1)].
(3) If $R$ is a left perfect ring (in the sense of [1, Theorem P]), then all flat left $R$-modules are projective, and therefore all left $R$-modules are cotorsion. Consequently, the class of $A/R$-flaprojective left $A$-modules coincides with the class of projective left $A$-modules in this case.
We will denote the class of all $R$-cotorsion left $A$-modules (i. e., left $A$-modules that are cotorsion over $R$) by $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}}\subset A{\operatorname{\mathsf{--Mod}}}$ and the class of all $A/R$-flaprojective left $A$-modules by $A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}\subset A{\operatorname{\mathsf{--Mod}}}$.
Given a left $R$-module $G$, the left $A$-module $A\otimes_{R}G$ is said to be induced from $G$. Notice that the $A$-module $A\otimes_{R}G$ is flat whenever the $R$-module $G$ is flat. The following lemma describes the Ext spaces from modules induced from flat modules.
Lemma 3.4.
Let $R\longrightarrow A$ be a homomorphism of associative rings, $G$ be a flat left $R$-module, and $M$ be a left $A$-module. Then there are natural isomorphisms of abelian groups $\operatorname{Ext}^{n}_{A}(A\otimes_{R}G,\>M)\simeq\operatorname{Ext}^{n}_{R}(G,M)$ for all integers $n\geq 1$.
Proof.
This is a particular case of [5, Proposition VI.4.1.3]. See also [10, formula (1.1) in Section 1.2] or [13, Lemma 1.1]. ∎
Theorem 3.5.
Let $R\longrightarrow A$ be a homomorphism of associative rings. Then
(a) the pair of classes of left $A$-modules $(\mathsf{F},\mathsf{C})$ $=$ $(A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}$, $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}})$ is a complete cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$;
(b) a left $A$-module is $A/R$-flaprojective if and only if it is a direct summand of an $A$-module filtered by the left $A$-modules $A\otimes_{R}G$ induced from flat left $R$-modules $G$.
Proof.
This is [12, Proposition 3.3]. Notice that the left $A$-module $A^{+}=\operatorname{Hom}_{\mathbb{Z}}(A,\mathbb{Q}/\mathbb{Z})$ is injective, hence pure-injective, hence cotorsion; so by Lemma 2.3,  $A^{+}$ is a cotorsion left $R$-module and [12, Proposition 3.3] is applicable.
To spell out some details, one starts with observing that the “if” assertion in part (b) holds by Lemmas 3.4 and 1.1. Let $\mathsf{S}_{0}$ be a set of flat left $R$-modules such that $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}=\mathsf{Fil}(\mathsf{S}_{0})$, as in Proposition 3.1. Then $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}=(\mathsf{S}_{0})^{\perp_{1}}\subset R{\operatorname{\mathsf{--Mod}}}$ by Lemma 1.1. Let $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$ be the set of all left $A$-modules $A\otimes_{R}S$, where $S\in\mathsf{S}_{0}$. Then it follows that $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}}=\mathsf{S}^{\perp_{1}}\subset A{\operatorname{\mathsf{--Mod}}}$ by Lemma 3.4. Without loss of generality we can assume that ${}_{R}R\in\mathsf{S}_{0}$; then ${}_{A}A\in\mathsf{S}$. Now part (a) is provided by Theorem 1.2(a) and the “only if” implication in part (b) by Theorem 1.2(b). ∎
We call the complete cotorsion pair $(\mathsf{F},\mathsf{C})$ $=$ $(A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}$, $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}})$ from Theorem 3.5 the flaprojective cotorsion pair or the $A/R$-flaprojective cotorsion pair in the category of left $A$-modules $A{\operatorname{\mathsf{--Mod}}}$.
Corollary 3.6.
Let $R\longrightarrow A$ be a homomorphism of associative rings. Then the following conditions are equivalent:
  1. (1)
    a left $A$-module is cotorsion whenever it is cotorsion as a module over $R$;
  2. (2)
    a left $A$-module is cotorsion if and only if it is cotorsion as a module over $R$;
  3. (3)
    all flat left $A$-modules are $A/R$-flaprojective;
  4. (4)
    the classes of flat left $A$-modules and $A/R$-flaprojective left $A$-modules coincide.
Proof.
(1) $\Longleftrightarrow$ (2) holds because all cotorsion $A$-modules are always cotorsion over $R$ by Lemma 2.3.
(3) $\Longleftrightarrow$ (4) holds because all $A/R$-flaprojective $A$-modules are always flat by Remark 3.3(1).
(2) $\Longleftrightarrow$ (4) The pair of classes of left $A$-modules ($A{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $A{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}$) is always a (complete) cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$ by the discussion in the beginning of this section. The pair of classes of left $A$-modules ($A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}$, $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}}$) is always a (complete) cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$ by Theorem 3.5(a).
In a cotorsion pair, any one of the two classes determines the other class as the respective right/left $\operatorname{Ext}^{1}_{A}$-orthogonal class in $A{\operatorname{\mathsf{--Mod}}}$. Therefore, one has $A{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}=A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}$ if and only if $A{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}=A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}}$. ∎
One can say that a ring homomorphism $R\longrightarrow A$ is relatively left perfect if any one of the equivalent conditions of Corollary 3.6 is satisfied. Theorem 2.8 above and following corollary claim that the ring homomorphism $R\longrightarrow A$ is relatively left perfect whenever $A$ is a finitely generated projective right $R$-module.
The following corollary is the second main result of this paper.
Corollary 3.7.
Let $R\longrightarrow A$ be a homomorphism of associative rings such that $A$ is a finitely generated projective right $R$-module. Then the classes of flat left $A$-modules and $A/R$-flaprojective left $A$-modules coincide. In other words, every flat left $A$-module is a direct summand of an $A$-module filtered by left $A$-modules $A\otimes_{R}G$ induced from flat left $R$-modules $G$.
Proof.
According to Corollary 3.6, the first assertion is an equivalent restatement of Theorem 2.8. The second assertion follows by Theorem 3.5(b). ∎
Finally, we can formally prove the theorem stated in the introduction.
Proof of Theorem 0.1.
Part (a) is Theorem 2.8. Part (b) is Corollary 3.7. ∎
Remark 3.8.
Let $R$ be a left perfect ring in the sense of [1, Theorem P], i. e., all flat left $R$-modules are projective. Let $R\longrightarrow A$ be a ring homomorphism making $A$ a finitely generated projective right $R$-module. Then it follows from Corollary 3.7 that all flat left $A$-modules are projective, too. So $A$ is a left perfect ring as well.
However, this assertion can be proved in a greater generality with the classical methods. The projectivity condition on the right $R$-module $A$ can be dropped here; it suffices to assume that $A$ is a finitely generated right $R$-module. Indeed, by [1, Theorem P(6)], a ring $S$ is left perfect if and only if all descending chains of cyclic right $S$-modules terminate. According to [4, Theorem 2], this holds if and only if all descending chains of finitely generated right $S$-modules terminate. It remains to observe that, whenever $A$ is a finitely generated right $R$-module, a right $A$-module is finitely generated if and only if it is finitely generated as an $R$-module.
Question 3.9.
Of course, neither Theorem 2.8 nor Corollary 3.7 are true without the condition that the right $R$-module $A$ is finitely generated. If this condition is dropped, it suffices to consider the case when $R=k$ is a field and $A=k[t]$ is the $k$-algebra of polynomials in one variable $t$. Then the $A$-module $k(t)$ of rational functions in $t$ and the $A$-module $k[t,t^{-1}]$ of Laurent polynomials in $t$ are flat but not projective. So there exist $A$-modules that are not cotorsion (in fact, the free $A$-module $A$ is not cotorsion). However, every $k$-vector space is a cotorsion $k$-module, so all $A$-modules are cotorsion over $k$. Therefore, all $A/k$-flaprojective $A$-modules are projective.
However, we do not know whether the assertions of Theorem 2.8 and/or Corollary 3.7 hold true with the projectivity assumption on the right $R$-module $A$ dropped. For example, if $A$ is an associative algebra over a commutative ring $R$ and the $R$-module $A$ is finitely presented, does it follow that the ring homomorphism $R\longrightarrow A$ is relatively (left) perfect?